OpenStudy (jmartinez638):

Write an equation for the tangent line to the graph of the following equation at the point (2,1):

OpenStudy (jmartinez638):

\[x^3 +y^3 = 4xy+1\]

OpenStudy (solomonzelman):

Did you find the derivative?

OpenStudy (jmartinez638):

I had a little trouble with that...

OpenStudy (solomonzelman):

What is the derivative of \(x^3\) ?

OpenStudy (jmartinez638):

3x^2

OpenStudy (solomonzelman):

yes, now how about y^3? (Apply the same rule, but multiply times y')

OpenStudy (jmartinez638):

I have a calculator that I can do all this on, but it ends up being a seriously detrimental crutch, so figuring out how to work it out is paramount. (<---- maybe some wordplay)

OpenStudy (jmartinez638):

3y^2 * y'

OpenStudy (solomonzelman):

yes, that is true

OpenStudy (solomonzelman):

Okay, now apply the product rule to 4xy

OpenStudy (solomonzelman):

what is the derivative of y? what is the derivative of 4x?

OpenStudy (jmartinez638):

It would be... 0? I feel like I am doing something wrong. Product rule is d/dx[f(x)g(x)]= f(x)g'(x) + g(x)f'(x)?

OpenStudy (jmartinez638):

Derivative of 4x would be four actually, not zero, because the integer just 'comes along for the ride' as it were.

OpenStudy (solomonzelman):

Yes, go on...

OpenStudy (jmartinez638):

And y would be 1 x y or y x y' or 1 x y'? A bit weird there...

OpenStudy (solomonzelman):

In general, the derivative of f(x) is f'(x) right?

OpenStudy (jmartinez638):

Yes, f'(x) = the derivative of x

OpenStudy (solomonzelman):

No, f'(x) is the derivative of the function f(x).

OpenStudy (solomonzelman):

right?

OpenStudy (solomonzelman):

In general, all I am asking is that you know that for any function f(x), the derivative is f'(x)... correct?

OpenStudy (jmartinez638):

Yes

OpenStudy (solomonzelman):

And now, you have to recall that y is also a function of x. So the derivative of y is y'

OpenStudy (jmartinez638):

Yes

OpenStudy (solomonzelman):

And this way, when you differentiate 4x•y, you get: (4)•y + 4x•(y')

OpenStudy (solomonzelman):

Am I correct?

OpenStudy (jmartinez638):

Sure, that makes sense.

OpenStudy (solomonzelman):

Ok, now, when you differeniate each component you get: \(\large\color{black}{ \displaystyle x^3 +y^3 = 4xy+1 }\) \(\large\color{black}{ \displaystyle 3x^2 +3y^2y' = 4y+4xy'+0 }\)

OpenStudy (anonymous):

May I ask are you finding the derivative with respect to y?

OpenStudy (solomonzelman):

No, dy/dx.

OpenStudy (anonymous):

Ok..watching and relearning

OpenStudy (solomonzelman):

It is a derivative with respect to x. Each y is a function of x, thus it deserves a chain rule of y'.

OpenStudy (solomonzelman):

jmartine Do you see how I get the equation below? \(\large\color{black}{ \displaystyle 3x^2 +3y^2y' = 4y+4xy'+0 }\)

OpenStudy (jmartinez638):

Yes, I do. "It is a derivative with respect to x. Each y is a function of x, thus it deserves a chain rule of y'." That makes a lot of sense

OpenStudy (solomonzelman):

Alright, now just using algebra you have to isolate the y'

OpenStudy (jmartinez638):

3x^2+3y^2y′=4y+4xy′+0 3x^2-4y=3y^2y' + 4xy' (3x^2-4y)/(3y^2 *4x)=2y' (3x^2-4y)/(3y^2 *4x*2)=y'

OpenStudy (jmartinez638):

So it would be \[y' = \frac{ 3x^2-4y }{ 3y^2 *4x*2}\]

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{3x^2 -4y}{{4x-3y^2}}=y' }\)

OpenStudy (jmartinez638):

Ok, I got a little messed up moving stuff from either side. I thought if you had 2x + 3x= 1 and you divided out 2 and 3 it would leave x + x or 2x? Maybe just me going blank on some core math stuff randomly...

OpenStudy (solomonzelman):

Well, just a matter of algebraic manipulation...

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y'=\frac{3x^2 -4y}{4x-3y^2} }\) gives you the expression for the function's slope. That means that at (2,1) the slope is: \(\large\color{black}{ \displaystyle y'=\frac{3(2)^2 -4(1)}{4(2)-3(1)^2}=\frac{12 -4}{8-3} =\frac{8}{5} }\)

OpenStudy (solomonzelman):

So your line becomes \(\large\color{black}{ \displaystyle y-y_1=m(x-x_1) }\) \(\large\color{black}{ \displaystyle y-1=\frac{8}{5}(x-2) }\)

OpenStudy (jmartinez638):

Ok, that definitely makes sense!

OpenStudy (solomonzelman):

:)

OpenStudy (jmartinez638):

Thanks for helping, I really understand how the whole process of solving the problem works now!